OpenStudy (anonymous):

1, 3, 4, 7, 8, 10 is this a geometric or an arithmetic series?

OpenStudy (anonymous):

What is the definition of a geometric and an arithmetic series?

OpenStudy (anonymous):

this is tough one, :)

OpenStudy (anonymous):

is it neither?

OpenStudy (anonymous):

that's what it looks like

OpenStudy (anonymous):

Neither.

OpenStudy (anonymous):

okay wht abt these othere ones? 0.3, 0.03, 0.003, ... -4, -12, -36, -108, ... -3, 6, -12, 24, -48, ... -1, 2, -4, -8, 16, ...

OpenStudy (anonymous):

Are you just posting homework?

OpenStudy (anonymous):

first one think of as \[\frac{3}{10},\frac{3}{100},\frac{3}{1000},...\]or as \[\frac{3}{10},\frac{3}{10^2},\frac{3}{10^3},...\] to see that it is geometric

OpenStudy (anonymous):

all geometric

OpenStudy (anonymous):

No wait. This sequence is interesting.

OpenStudy (anonymous):

Phi-antipalindromic numbers

OpenStudy (anonymous):

http://www.youtube.com/watch?v=XZPpG9qPqgA

OpenStudy (anonymous):

How did you realize that? :D

OpenStudy (anonymous):

When it comes to numbers I know a few things ;)

OpenStudy (anonymous):

im making sure my answers are right @noliec. I dont ask ppl to do my homework thers no benefit in it for me.

OpenStudy (anonymous):

Then I apoligize gr12! :)

OpenStudy (anonymous):

i have another question: If a geometric sequence has t4 = 3 and t8 = 243, find t11.

OpenStudy (anonymous):

is term 11 positive or negative? or is it both?

OpenStudy (anonymous):

and what's your answer to check? :)

OpenStudy (anonymous):

my answer is 6561, but im unsure of if its positive or neagtive? cause you do to the fouth rot

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I believe it's +ve.

OpenStudy (anonymous):

yea thts wht i thought. one last question please? i absolutely have no idea how to solve this? If 6, 6^x + 2, and 6^3x - 3 are the first three terms of a geometric sequence, what is the value of x?

OpenStudy (anonymous):

if a, b and c are in geometric progression, what can we say for sure?

OpenStudy (anonymous):

Btw it could be\( \pm \) for your other question.

OpenStudy (anonymous):

okay, and wht abt the question i just asked.